. In a population of 215 frogs, 140 are green, 50 are brownish green, and 25 are brown. The allele for green is denoted GG, while that for brown is GB, and these two alleles show incomplete dominance relative to each other.

a. What are the numbers of GG and GB alleles in this group of frogs?

b. What are the allele frequencies of GG and GB in this population?

c. What are the expected frequencies of the phenotypes (green, brownish green, brown) if the population were at Hardy-Weinberg equilibrium.

d. Based on these expected frequencies, what would the original numbers of green, brownish green and brown individuals EXPECTED to be were they in Hardy-Weinberg equilibrium? In other words are the numbers 140 green, 50 brownish green, and 25 brown individuals in Hardy-Weinberg equilibrium given the expected ratios calculated in part c?

e. Do the OBSERVED numbers of 140 green, 50 brownish green, and 25 brown individuals in the population differ SIGNIFICANTLY from the EXPECTED numbers calculated in Part d.? Looks like time for a X2 analysis.

f. Could a case be made that there was considerable inbreeding in this population of frogs? Explain.
Answer #1

Question-1: a. What are the numbers of GG and GB alleles in this group of frogs?
The probability of GG is, 215/405 = 0.53
The probability of GB is, 25/405 = 0.061

Question-2: What are the allele frequencies of GG and GB in this population?
Let us assume that the gene coding for greenish brown is BG, they are heterozygous
The frequency of BG is = 50/405 = 0.123
The allele frequency of GG is = the frequency of GG + ½ (frequency of BG)
= 0.53 + ½ (0.123) = 0.53 + 0.0615 = 0.591
The allele frequency of GB is = the frequency of GB + ½ (frequency of BG)

The allele frequency of GB is = 0.061+ 0.0615 = 0.1225